OpenStudy (anonymous):

What is the exact value of the expressions the square root of 180. − the square root 125. + the square root of 5.? Simplify if possible.

OpenStudy (anonymous):

\[\sqrt{180}-\sqrt{125}+\sqrt{5}\]